OpenStudy (anonymous):

Did I get the right answer? if f(x)=x^2+3x+5 what is f(a+h)? A. a^2+2ah+h^2+3a+3h+5 B. (a+h)^2+3(a+h)+5(a+h) C. (x^2+3ax+5) (a+h) D. h^2+3a+3h+5 Is it C? I'm not exactly sure what the question is asking me.

OpenStudy (anonymous):

subtitute x with a+h

OpenStudy (anonymous):

@febylailani So the answer will be B then?

OpenStudy (anonymous):

no, why B? there's no x after 5 so you mustn't put a+h after 5. do you think so?

OpenStudy (anonymous):

f(x)=x^2+3x+5 f(a+h)=(a+h)^2 +3(a+h)+5 :)